Prev Next

SW&IT Meeting Minutes 2021-11-08

Attendees: Brandon, David, Glenn, Matthieu, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20211101
LR20211101 - Bec, Matthieu 40.0
LR20211101 - Mechtley, Brandon 0.0
LR20211101 - Eychaner, Glenn 0.0
LR20211101 - Edmeades, David 40.0
LR20211101 - Grenz, Paul 20.0
LR20211101 - Hooper, Stephen 0.0
LR20211101 - Zhang, Yang 40.0
LR20211101 - Zhang, Xianyu 40.0

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

  • sciops new domain names requests - lbtoscience.org or sciencelbto.org, not both? maybe they need some guidance
  • DPM winter time @3pm (Pat updated invite)

Team reports

MB

  • Last week
    • OC, Management
    • Performance Appraisals
    • ATT PM#10
    • learning some more about AO s/w, join discussions and scheming on #ao-sw
    • TCS/AOS/RTDB string bug
    • LBTplot conversion to DMS (dimm/guide)
  • This week
    • Performance Appraisals
    • git/github "policy" (= philosophy slides) for new staff
    • LBTplot conversion to DMS (ctd.)

DE

  • Infra2 work, very dead.
    • Tech visit this morning; not sure of exact timing but I'll probably not be here for the meeting
    • Replacing motherboard, RAM, and CPU.
  • Backup server finally got Windows 21H1 installed, will get the backup service going this week
  • Options for escalating Windows Server?
  • Re-order SSD for Synology
  • Followup on new Synology orders
  • New UPS battery for mask cutter machine, sent to RR

YZ

  • Previous:
    • Some LBC stuff (lbc-alerts list reconfiguration).
    • A new AOS patch released (to be activated next eng night on Nov 14).
    • Some TMS work (and discussion with Andrew on some remaining issues in TMS algorithms).
  • Next:
    • Some LBC and MODS issues (e.g., WCS, several issues to discuss with Olga and others, IT8472, IT8509 etc).
    • Routine TMS work.
    • Continue testing and debugging with Paul on OSS/Argos swingarm (and maybe OSS patch).

PSG

  • Last week
    • Updated the HL2000 (white light source) controller.
    • Tested controller such that it can maintain multiple connections.
  • This week
    • Final OSS testing to show status is correct for ARGOS swing arms (Depoying vs Retracting)
    • Run through white light source one final time before release.
    • Run through swing arms one more time before release.
    • Get virtual machine configured.

GE

  • Last week
    • Digging through the 550,000 lines of code of the LUCI system.
    • Created a test version of LUCI software that fixes the image-cube issue and updates Aladin to version 11.
    • Found that the RTD software does not obey the restrictions of only modifying the GUI from the Event Dispatch Thread.
      • I hope this problem does not extend beyond the RTD.
    • Found that Aladin expects commands to come from a single thread which is not the Event Dispatch thread, while the RTD sends commands from any of several RMI threads and the Event Dispatch thread. This expectation is not documented explicitly anywhere I can find, but is implicit in the design of cds/aladin/Command.java and the surrounding synchronization.
  • This week
    • Digging through the 550,000 lines of code of the LUCI system.
    • Continue working on the RTD issues.
    • Examine the other LUCI GUIs to check for GUI modifications from threads other than the Event Dispatch Thread

BM

  • Last week
    • AO code deployment plan
    • Small fixes to gain ramp IDL
    • Proposed DiagnApp /Housekeeper changes. Waiting for Alfio to review. (maybe just PR?)
    • New issues (most from LBTI run)
      • 2x PauseAo in a row clear gains file (?--might want to duplicate bug)
      • ResumeAo after manual LoadGainTag overwrites new gains (discuss at AOG Wednesday)
      • SkipFrames (JCG interested in state of actuators before StopAo)
      • Python app names not preserved in logs
  • This week:
    • M: Discuss AO code deployment, SW testing
    • W: Power backplane meeting w/MG
    • W: AOG
    • DiagnApp /HouseKeeper changes (logging during partial timeouts)
    • Look into/fix issues from LBTI run.

SH

  • Ignition (SCADA) VMs for Dan Rapoza
  • Monitors for remote observing
  • Omada controller on mountain
  • MCSPU stuff (maybe obs1 stuff) on Saturday morning.
  • VPN stuff
  • Cisco quote ping.
  • Mark Smithwright's computer.

next:

  • Fedora list, Fedora discussion
  • More orders
  • Training on Wednesday morning
  • Thursday is a a U. Holiday (Veteran's day).
  • obs5 and obs6
Topic revision: r9 - 08 Nov 2021, MatthieuBec
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback